# Resource Allocation for Twin Maintenance and Computing Task Processing in Digital Twin Vehicular Edge Computing Network

### v1

@@ Abstract
As a promising technology, vehicular edge computing (VEC) can provide computing and caching services by deploying VEC servers near vehicles. However, VEC networks still face challenges such as high vehicle mobility. Digital twin (DT), an emerging technology, can predict, estimate, and analyze real-time states by digitally modeling objects in the physical world. By integrating DT with VEC, a virtual vehicle DT can be created in the VEC server to monitor the real-time operating status of vehicles. However, maintaining the vehicle DT model requires ongoing attention from the VEC server, which also needs to offer computing services for the vehicles. Therefore, effective allocation and scheduling of VEC server resources are crucial. This study focuses on a general VEC network with a single VEC service and multiple vehicles, examining the two types of delays caused by twin maintenance and computational processing within the network. By transforming the problem using satisfaction functions, we propose an optimization problem aimed at maximizing each vehicle’s resource utility to determine the optimal resource allocation strategy. Given the non-convex nature of the issue, we employ multi-agent Markov decision processes to reformulate the problem. Subsequently, we propose the twin maintenance and computing task processing resource collaborative scheduling (MADRL-CSTC) algorithm, which leverages multi-agent deep reinforcement learning. Through experimental comparisons with alternative algorithms, it demonstrates that our proposed approach is effective in terms of resource allocation.
@@
Digital twin, Twin maintenance,Vehicular edge computing, Resource allocation
@@ I Introduction
The combination of DT and VEC can not only collect real-time operational data of vehicles in the VEC network, but also carry out the real time control and change of the status of vehicles[41]. Real-time vehicle information can be accessed and compared with historical data to identify and address potential problems or risks in vehicle operation proactively[42]. However, this integration also presents challenges, particularly in terms of resource management. Establishing and maintaining DT models requires computing resources for information synchronization, while processing vehicle computing tasks on the server further consumes resources. Therefore, it is crucial to explore how to develop reasonable resource allocation strategies to prevent delays in task execution.
@@
In this paper, we consider a DT scenario for mobile edge network with multiple vehicles associated with a single VEC server and propose a multi-agent based twin maintenance and computing task processing resource collaborative scheduling algorithm (MADRL-CSTC) [^1] The source code has been released at: https://github.com/qiongwu86/Resource-allocation-for-twin-maintenance-and-computing-tasks-in-digital-twin-mobile-edge-network., which maximizes the resource utility of computing resource allocation while ensuring the time limit requirements of twin maintenance and vehicle computing tasks. The main contributions of this article are summarized as follows:
  • We consider a general scenario of a DT mobile edge network with multiple vehicles and a single VEC server. We first analyzed the two types of delays caused by simultaneous vehicle digital twin maintenance and computational task processing in this scenario. Then, describe these two types of delays as the resource utility under the current server computing resource allocation strategy using a satisfaction function. Afterwards, considering various constraints, the optimization problem of maximizing the resource utility of each vehicle is proposed.
  • Due to the non convexity of the optimization problem, taking into account the total server resources in the scenario, the deadline for twin maintenance and computing tasks, the maximum transmission power of vehicles, interference between vehicles and servers, and vehicle mobility, we reconstructed the problem using a multi-agent Markov decision process (MDP). Specifically, each vehicle is modeled as an intelligent agent interacting with the environment, i.e. other intelligent agents. Considering the problem itself, we set the reward to the resource utility of each vehicle.
  • In order to explore the optimal allocation strategy for maximizing the utility of vehicle resources, we propose the MADRL-CSTC algorithm. Specifically, we utilized centralized training, distributed execution, and actor-critic (AC) framework, while designing estimation and target AC neural networks for policy iteration and value iteration, respectively. We also use $\epsilon$-greedy algorithm to balance the relationship between exploration and utilization. By comparing it with other algorithms through extensive experiments, the superiority and effectiveness of our algorithm can be verified.
@@
The remaining part of this article is organized as follows. Section II reviews the related work. Section III introduces the system model and analyzes two types of delays caused by twin maintenance and computational task processing, and then uses satisfaction functions to formulate optimization problems. In section IV, we reconstruct the problem using multi-agent MDP and propose the MADRL-CSTC algorithm. Section V presents simulation results to demonstrate the effectiveness of our proposed algorithm. Section VI draws conclusions.
@@ II-B Digital Twin Vehicular Edge Computing
Recently, there have been some studies combining DTs with VEC. In [33], Dai et al. proposed an adaptive DT VEC network that utilizes DT caching content and designs an offloading scheme based on the DRL framework to minimize offloading latency. In [48], Zhang et al. combined DTs with artificial intelligence into the VEC network, guiding vehicles to aggregate edge services by deploying digital twins in road side units (RSUs) to minimize offloading costs. In [49], Liao et al. considered that in mixed traffic scenarios, the planned operation of autonomous vehicles (CAVs) may be affected by human driven vehicles (HDVs), and developed a driver DT model that was deployed on the server to help CAVs predict the possible lane changing behavior of surrounding HDVs. In [50], Zhang et al. proposed a socially aware vehicle edge caching mechanism, which constructs a vehicle social relationship model by deploying a digital twin model of edge caching systems on RSUs and designs an optimal caching scheme based on DRL. In [51], Zheng et al. proposed that vehicles in heterogeneous vehicle networks need to improve communication efficiency by selecting different networks, while also facing situations of uneven vehicle distribution and dynamic networks. They established a prediction model in DT to predict the waiting time for vehicles to connect to the network and return the results to the vehicles for decision-making.In [52], Zhao et al. established a vehicle DT model in RSUs to learn the global information of the vehicle, thereby assisting clustering algorithms in reducing the scope of vehicle task offloading and achieving the goal of offloading prediction. In [53], Sun et al. proposed the Digital Twin Edge Network (DITEN) and established DT models for edge servers and the entire MEC system to provide auxiliary services for mobile user business offloading, thereby minimizing offloading latency. In [54], Li et al. integrated unmanned aerial vehicles (UAVs) into VEC and established DT models for RSUs, UAVs, and vehicles in the central controller to manage UAV and RSU resources and assist in offloading vehicular tasks.Although establishing a DT model of vehicles in the servers of the VEC network can help vehicles perform better task offloading, edge business aggregation, or execute specific traffic behaviors, at the same time, the maintenance of the vehicle DT model will also consume server resources, which also needs to be considered.
As mentioned above, there has been no research on resource allocation issues in the coexistence of vehicle twin maintenance and VEC server services for vehicles.
@@ III System model
As shown in Figure 1, we consider a DT scenario of Mobile Edge Network (DTMEN) with $N$ vehicles and a server equipped base station. For clarity, let $N\stackrel{{\scriptstyle\Delta}}{{=}}\{1,2,3,...,N\}$ represent the number of vehicles driving on the lane, and its initial position follows the Poisson clustering process. Specifically, these vehicles will travel in one direction along the lane and generate computing tasks during their journey. Due to their limited computing resources, the vehicles will connect to the base station (BS) through the V2I link using a cellular interface (Uu) to offload the computing tasks to the server for processing. The completed results will be returned to the vehicles. In addition, DT models are established for these vehicles on the server to obtain real-time operational status of the vehicles. Meanwhile, in order to maintain the digital twin models of these vehicles, the server needs to synchronize information with the vehicles to ensure that the models can accurately reflect the driving conditions of the vehicles.
III-A Vehicle Movement Model and Channel Model
III-A1 Vehicle Movement
Similar to [55], a spatial orthogonal coordinate system is established with the base station as the origin in figure 2. The positive direction of the x-axis is the direction of the vehicle traveling along the lane, that is, east, the positive direction of the y-axis is south, and the positive direction of the z-axis is along the direction of the base station antenna. Therefore, the coordinate of vehicle $i(i\in\{1,2,3...,N\})$ located on lane j at time slot t can be represented as $PO_{i}(t)$
$PO_{i}(t)=(X_{i,j}(t),Y_{i,j}(t),0),i\in\{1,2,3...,N\},$ (1)
where $X_{i,j}(t)$ is the horizontal ordinate of vehicle $i$ driving on lane $j$ in time slot $t$, $Y_{i,j}(t)$ is the ordinate of vehicle $i$.
@@ III-A2 Channel Model
where $v_{i,B}^{t}$ is the large-scale fading component between vehicle $i$ and BS, which is composed of path loss and shadow distribution. $h_{i,B}^{t}$ is the small-scale path fading component between vehicle $i$ and BS at time slot $t$, which follows a circularly symmetric complex Gaussian distribution with unit variance. We use a first-order Gaussian Markov process to describe small-scale path fading. Therefore, the update of $h_{i,B}^{t}$ can be represented as
$h_{i,B}^{t}=\kappa h_{i,B}^{t-1}+\frac{1}{(\sqrt{(PO_{i}(t)-P_{B}(t))^{2}})^{\beta}},$ (5)
where $\kappa$ is correlation coefficient, $P_{B}(t))^{2}$ is the coordinates of the base station, $\beta$ is a path loss index.
III-B Communication Model and Computing Model
III-B1 Communication Model
As shown in Figure 1, the communication between the vehicle and the base station includes information updates for vehicle twin maintenance, uplink communication for task transmission with the same server, and downlink communication for returning calculation results to the vehicle. It should be noted that since the size of the results processed by the digital twin edge server is much smaller than before, and the communication rate of the downlink is much larger, Therefore, we ignore the delay caused by downlink communication. According to Shannon’s theorem, we can calculate the uplink communication rate between the vehicle and the base station as follows[60]
@@ III-B2 Computing Model
After establishing the digital twin of vehicle $i$, in order for the digital twin model to accurately reflect the current situation of the vehicle, the digital twin of vehicle $i$ in the VEC server needs to be synchronized with the information of vehicle $i$ in the physical space[38]. We let $\Gamma_{i}(t)=\{{D_{i}^{dt}(t),C_{i}^{dt}(t),T_{i}^{dt}}\}$ represents the information of vehicle $i$ that needs to be synchronously updated at the current time slot $t$, where $D_{i}^{dt}(t)$ is the size of the information of vehicle $i$ that needs to be updated at the current time slot $t$, $C_{i}^{dt}(t)$ represents the CPU frequency required to update the unit size of vehicle $i$ information, and $T_{i}^{dt}$ is the maximum delay limit for updating vehicle $i$ information. The updated information needs to be transmitted from vehicle $i$ to the server equipped with the base station first, and then the server will allocate computing resources to maintain the digital twin of vehicle $i$. Here, according to the hosting restrictions of [61], the maintenance of digital twins on vehicle $i$ is also related to the total number of digital twins it needs to maintain on its server. In addition, we record the total amount of computing resources on the server as $F$, and the computing resources requested by vehicle $i$ for twin maintenance as $f_{i}^{dt}$. Therefore, the time $t_{i}^{dt}$ required to complete the digital twin maintenance of vehicle $i$ can be calculated as
$t_{i}^{dt}=\frac{D_{i}^{dt}(t)}{r_{i,B}^{t}}+\frac{D_{i}^{dt}(t)C_{i}^{dt}(t)}{f_{i}^{dt}}N,$ (7)
@@
Vehicles may generate computing tasks while driving, and due to their insufficient computing power, they may choose to offload the computing tasks to edge servers for processing. We represent the computing task generated by vehicle $i$ at time slot $t$ as $\zeta_{i}(t)=\{{D_{i}^{tk}(t),C_{i}^{tk}(t),T_{i}^{tk}}\}$, where $D_{i}^{tk}(t)$ is the computing task size generated by vehicle $i$ at time slot $t$, $C_{i}^{tk}(t)$ is the CPU frequency required to compute unit size tasks, and $T_{i}^{tk}$ is the maximum delay limit for processing computing tasks. Similar to the previous maintenance of digital twins, the vehicle $i$ will first offload the computing task to the server, and then the server will allocate computing resources to handle the computing task. Therefore, the time $t_{i}^{tk}$ required to complete the calculation task, which can be calculated as
$t_{i}^{tk}=\frac{D_{i}^{tk}(t)}{r_{i,B}^{t}}+\frac{D_{i}^{tk}(t)C_{i}^{tk}(t)}{f_{i}^{tk}},$ (8)
@@
where $f_{i}^{tk}$ represents the computing resources requested by vehicle $i$ for processing computing tasks transmitted to the server.
@@
It should be noted that the maintenance and computing tasks of the digital twin of vehicle $i$ are carried out simultaneously, so the sum of the computing resources $f_{i}^{dt}$ allocated for the maintenance of the digital twin of vehicle $i$ and the computing resources $f_{i}^{tk}$ allocated for processing the computing tasks generated by vehicle $i$ can’t exceed the total computing resources $F$ of the server, that is, $\Sigma_{1}^{N}f_{i}^{dt}+f_{i}^{tk}\leq F$.
III-C Optimization Problem

### v2

# Resource Allocation for Twin Maintenance and Task Processing in Vehicular Edge Computing Network

@@ Abstract
In the digital twin mobile edge network, the maintenance of the vehicle twin model and vehicular task processing in the server require the support of computing resources. In addition, they are performed simultaneously. Therefore, how to allocate resources for twin maintenance and task processing under limited server resources is crucial. However, current research tends to ignore the aspect of resource competition for twin maintenance. In this study, we analyze the delays of these two affected by resource allocation under a generic digital twin mobile edge network (DTMEN) to construct the optimization problem. For this problem, we transformed the problem using a Markov decision process. Meanwhile, we propose a multi-agent reinforcement learning (MADRL) based twin maintenance and task processing resource collaborative scheduling (TMTPRCS) algorithm to solve the problem. Experiments show that our proposed approach is effective in terms of resource allocation compared to other alternative algorithms.
@@
Twin maintenance, Vehicular edge computing, Resource allocation
@@ I Introduction
The combination of DT and VEC can not only collect real-time operational data of vehicles in the VEC network, but also carry out the real time control and change of the status of vehicles[34]. Real-time vehicle information can be accessed and compared with historical data to identify and address potential problems or risks in vehicle operation proactively[35]. By deploying the DT at the VEC server, the DT can continuously monitor the physical state of the vehicle and network environment. This allows proactive decision making based on the latest data. In addition, DT acts as a virtual representation of the physical system, allowing decisions to be tested and validated in a risk-free environment, as opposed to deploying DRL directly to the server. This enhances the robustness of the decision-making process. However, this integration also presents challenges, particularly in terms of resource management. Current research under DTMEN tends to ignore the fact that digital twin maintenance can place demands on computing resources. Under DTMEN, vehicle digital twins are often built in VEC server. In addition, the VEC server will likewise need to allocate resources for vehicular computing tasks. And they are often performed simultaneously. Therefore, it becomes critical to allocate resources to meet the needs of both.
@@
In this paper, we consider a DT scenario for mobile edge network with multiple vehicles associated with a single VEC server and propose a multi-agent based twin maintenance and computing task processing resource collaborative scheduling algorithm (TMRCS) [^1] The source code has been released at: https://github.com/qiongwu86/Resource-allocation-for-twin-maintenance-and-computing-tasks-in-digital-twin-mobile-edge-network., which maximizes the resource utility of computing resource allocation while ensuring the time limit requirements of twin maintenance and vehicle computing tasks. The main contributions of this article are summarized as follows:
  • We consider a general scenario of a DT mobile edge network with multiple vehicles and a single VEC server. We analyze two types of delays under this network that are caused by the resource allocation policy.
  • In this paper, we focus on the problem of resource allocation when twin maintenance is performed simultaneously with on-board tasks, which has been neglected in the existing literature. Therefore, we characterize the two types of delays using the satisfaction function and construct the associated optimization problem taking into account various constraints.
  • We propose a MADRL-based TMTPRCS algorithm to solve the optimization problem. The effectiveness and superiority of our proposed method is verified by comparing it with other algorithms.
@@
The remaining part of this article is organized as follows. Section II reviews the related work. Section III introduces the system model and analyzes two types of delays caused by twin maintenance and computing task processing, and then uses satisfaction functions to formulate optimization problems. In section IV, we reconstruct the problem using multi-agent MDP and propose a corresponding algorithm. Section V presents simulation results to demonstrate the effectiveness of our proposed algorithm. Section VI draws conclusions.
@@ II-B Digital Twin Vehicular Edge Computing
Recently, there have been some studies combining DTs with VEC. In [26], Dai et al. proposed an adaptive DT VEC network that utilizes DT caching content and designs an offloading scheme based on the DRL framework to minimize offloading latency. In [41], Zhang et al. combined DTs with artificial intelligence into the VEC network, guiding vehicles to aggregate edge services by deploying digital twins in road side units (RSUs) to minimize offloading costs. In [42], Liao et al. considered that in mixed traffic scenarios, the planned operation of autonomous vehicles (CAVs) may be affected by human driven vehicles (HDVs), and developed a driver DT model that was deployed on the server to help CAVs predict the possible lane changing behavior of surrounding HDVs. In [43], Zhang et al. proposed a socially aware vehicle edge caching mechanism, which constructs a vehicle social relationship model by deploying a digital twin model of edge caching systems on RSUs and designs an optimal caching scheme based on DRL. In [44], Zheng et al. proposed that vehicles in heterogeneous vehicle networks need to improve communication efficiency by selecting different networks, while also facing situations of uneven vehicle distribution and dynamic networks. They established a prediction model in DT to predict the waiting time for vehicles to connect to the network and return the results to the vehicles for decision-making.In [45], Zhao et al. established a vehicle DT model in RSUs to learn the global information of the vehicle, thereby assisting clustering algorithms in reducing the scope of vehicle task offloading and achieving the goal of offloading prediction. In [46], Sun et al. proposed the Digital Twin Edge Network (DITEN) and established DT models for edge servers and the entire MEC system to provide auxiliary services for mobile user business offloading, thereby minimizing offloading latency. In [47], Li et al. integrated unmanned aerial vehicles (UAVs) into VEC and established DT models for RSUs, UAVs, and vehicles in the central controller to manage UAV and RSU resources and assist in offloading vehicular tasks. Although establishing DT models of vehicles in the server of the VEC network can help vehicles perform better task offloading, edge business aggregation, or execute specific traffic behaviors, at the same time, the maintenance of each vehicle’s DT model will also consume server resources, which also needs to be considered.
As mentioned above, most of the literature related to digital twins does not take into account the fact that the maintenance of the twin model consumes server computing resources.
@@ III System model
As shown in Figure 1, we consider a DT scenario of Mobile Edge Network with $N$ vehicles and a server equipped base station. For clarity, let $\mathbf{N}\stackrel{{\scriptstyle\Delta}}{{=}}\{1,2,3,...,N\}$ represent the number of vehicles driving on the lane, and its initial position follows the Poisson clustering process. Specifically, these vehicles will travel in one direction along the lane and generate computing tasks during their journey. Due to their limited computing resources, the vehicles will connect to the base station (BS) through the V2I link using a cellular interface to offload the computing tasks to the server for processing. Meanwhile, in order to maintain the digital twin models of these vehicles, the server needs to synchronize information with the vehicles to ensure that the models can accurately reflect the driving conditions of the vehicles. It is noted that we assume that vehicles are always traveling within the coverage area of the current base station. Therefore the migration of the vehicle DT model is not considered.
III-A Digital Twin Model
The edge server maintains the DT model of vehicle $i$ and the DT model is denoted as
$DT_{i}=\{conf_{i},His_{i},\Gamma_{i}(t)\},$ (1)
where $conf_{i}$ is configuration information for vehicle $i$, $His_{i}$ is historical information of vehicle $i$, $\Gamma_{i}(t)$ is the information about the operation of vehicle $i$ in time slot $t$, which can be represented as
$\Gamma_{i}(t)=\{{D_{i}^{\text{dt}}(t),C_{i}^{\text{dt}}(t),T_{i}^{\text{dt}}}\},$ (2)
where $D_{i}^{\text{dt}}(t)$ is the size of the information, $C_{i}^{\text{dt}}(t)$ represents the CPU frequency required to update the unit size of vehicle $i$ information, and $T_{i}^{\text{dt}}$ is the maximum delay limit. Here, the configuration information $conf_{i}$ mainly refers to the parameters of the vehicle’s own hardware, which are fixed. Historical information $His_{i}$, on the other hand, mainly consists of data such as the vehicle’s historical mileage as well as historical fuel consumption. Therefore, these two symbols are not related to $\Gamma_{i}(t)$.
Note that the DT model developed in this paper is simple. In practical situations, the DT model will not only reflect physical entities, but will likewise reflect network topology relationships. In addition, it has functions such as virtual prediction as well as feedback control.
III-B Vehicle Movement Model and Channel Model
III-B1 Vehicle Movement
Similar to [48], a spatial orthogonal coordinate system is established with the base station as the origin in figure 2. This linear movement model is a simulation of the ideal driving condition of a vehicle on a highway with fewer cars. The real situation is much more complex. The positive direction of the x-axis is the direction of the vehicle traveling along the lane, that is, east, the positive direction of the y-axis is south, and the positive direction of the z-axis is along the direction of the base station antenna. Therefore, the coordinate of vehicle $i\in\mathbf{N}$ located on lane j at time slot t can be represented as $PO_{i}(t)$
$PO_{i}(t)=(X_{i,j}(t),Y_{i,j}(t),Z_{i,j}(t)),i\in\{1,2,3...,N\},$ (3)
where $X_{i,j}(t)$ is the horizontal ordinate of vehicle $i$ driving on lane $j$ in time slot $t$, $Y_{i,j}(t)$ is the ordinate of vehicle $i$, $Z_{i,j}(t)$ is the vertical coordinate of the vehicle, which we consider as 0.
@@ III-B2 Channel Model
where $v_{i,B}^{t}$ is the large-scale fading component between vehicle $i$ and BS, which is composed of path loss and shadow distribution, where path loss is calculated as $128.1+37.6\log_{10}(dis(v_{i},BS))$ and shadowing obeys the log-normal distribution. $dis(v_{i},\text{BS})$ is the distance between vehicle $i$ and BS. $h_{i,B}^{t}$ is the small-scale path fading component between vehicle $i$ and BS at time slot $t$, which follows a circularly symmetric complex Gaussian distribution with unit variance, i.e $h_{i,B}\sim\mathcal{CN}(0,1)$. We use a first-order Gaussian Markov process to describe small-scale path fading. Therefore, the update of $h_{i,B}^{t}$ can be represented as
$h_{i,B}^{t}=\kappa h_{i,B}^{t-1}+e_{i,B}^{t},$ (7)
where $e_{i,B}^{t}$ is the channel innovation process with distribution $\mathcal{CN}(0,1-\kappa^{2})$. the correlation coefficient $\kappa$ is denoted as $\kappa=J_{0}(2\pi f_{d}\tau)$, where $J_{0}(\bullet)$ is the zeroth order Bessel function. $f_{d}$ is the maximum Doppler frequency, which is calculated as $f_{d}=\frac{v_{i}f_{c}}{c}$, where $f_{c}$ is the carrier frequency, $c$ is the transmission rate of electromagnetic waves.
III-C Communication Model and Computing Model
III-C1 Communication Model
As shown in Figure 1, the communication between the vehicle and the base station includes information updates for vehicle twin maintenance, uplink communication for task transmission with the same server, and downlink communication for returning calculation results to the vehicle. It should be noted that since the size of the results processed by the digital twin edge server is much smaller than before, and the communication rate of the downlink is much larger. Therefore, we ignore the delay caused by downlink communication. According to Shannon’s theorem, we can calculate the uplink communication rate between the vehicle and the base station as follows[53]
@@ III-C2 Computing Model
After the DT model is created in the VEC server, DT model maintenance is important. Digital twin maintenance refers to the process of synchronizing information between the twin model and its physical entity in real time, i.e., the digital twin model self-updates itself by profitably processing real-time information transmitted by its physical entity. In this paper, we mainly focus on the process of processing information by the twin model. The updated information needs to be transmitted from vehicle $i$ to the server equipped with the base station first, and then send a request signal to its DT model for requesting a portion of the server resources to maintain the model. Here, according to the colocation restrictions of [54], the maintenance of digital twins on vehicle $i$ is also related to the total number of digital twins it needs to maintain on its server. In addition, we record the total amount of computing resources on the server as $F$, and the computing resources requested by vehicle $i$ for digital twin maintenance as $f_{i}^{\text{dt}}$. Therefore, the time $t_{i}^{\text{dt}}$ required to complete the digital twin maintenance of vehicle $i$ can be calculated as
$t_{i}^{\text{dt}}=\frac{D_{i}^{\text{dt}}(t)}{r_{i,B}^{t}}+\frac{ND_{i}^{\text{dt}}(t)C_{i}^{\text{dt}}(t)}{f_{i}^{\text{dt}}},$ (9)
Digital twin maintenance is modeled here as the process by which the data used by vehicles to maintain their models is synchronized to the digital models in the server. At the same time the process is influenced by the total number of twin models that the server has to maintain. This process is embodied in the algorithm designed in Section IV in that it acts on the rewards at each time step of the intelligence, which in turn directly affects policy learning.
@@
Vehicles may generate computing tasks while driving, and due to their insufficient computing power, they may choose to offload the computing tasks to edge servers for processing. We represent the computing task generated by vehicle $i$ at time slot $t$ as $\zeta_{i}(t)=\{{D_{i}^{\text{tk}}(t),C_{i}^{\text{tk}}(t),T_{i}^{\text{tk}}}\}$, where $D_{i}^{\text{tk}}(t)$ is the computing task size generated by vehicle $i$ at time slot $t$, $C_{i}^{\text{tk}}(t)$ is the CPU frequency required to compute unit size tasks, and $T_{i}^{\text{tk}}$ is the maximum delay limit for processing computing tasks. Similar to the previous maintenance of digital twins, the vehicle $i$ will first offload the computing task to the server, and then request computing resources to handle the computing task. Therefore, the time $t_{i}^{\text{tk}}$ required to complete the calculation task, which can be calculated as
$t_{i}^{\text{tk}}=\frac{D_{i}^{\text{tk}}(t)}{r_{i,B}^{t}}+\frac{D_{i}^{\text{tk}}(t)C_{i}^{\text{tk}}(t)}{f_{i}^{\text{tk}}},$ (10)
@@
where $f_{i}^{\text{tk}}$ represents the computing resources requested by vehicle $i$ for processing computing tasks transmitted to the server.
@@
It should be noted that the maintenance and computing tasks of the digital twin of vehicle $i$ are carried out simultaneously, so the sum of the computing resources used for twin maintenance and computing tasks processing for all vehicles can’t exceed the total computing resources $F$ of the server, that is, $\Sigma_{1}^{N}(f_{i}^{\text{dt}}+f_{i}^{\text{tk}})\leq F$.
III-D Optimization Problem